OpenStudy (anonymous):

for the pair of functions f(x)=^2+4 and g(x)=x+5, find the following a).(f*g)(x) b).(f*g)(2) c).(g*f)(x) d).(g*f)(2)

myininaya (myininaya):

f(x)=x^2+4?

myininaya (myininaya):

*<- this is multiplication?

OpenStudy (anonymous):

When you're multiplying functions, you just multiply whatever the function is equal to. \[(f*g)(x)=(x^2+2)(x+5)\]You just need to foil that to make it work. Make sure that you're not actually supposed to be doing a COMPOSITION of functions (see the attachment)

myininaya (myininaya):

(f*g)(x)=f(x)*g(x) or do you mean (f*g)(x)=f(g(x))

myininaya (myininaya):

* what is the notation you want this to mean?

OpenStudy (anonymous):

I think i mean f(g(x))

myininaya (myininaya):

okay so we will let this be the notation for composition functions (f*g)(x)=f(g(x))

myininaya (myininaya):

* eventhough we usually mean to this for multiplication

myininaya (myininaya):

\[(g*f)(x)=g(f(x))=g(x^2+4)=(x^2+4)+5=x^2+9\] \[(f*g)(x)=f(g(x))=f(x+5)=(x+5)^2+4=x^2+10x+25+4=x^2+10x+29\]

myininaya (myininaya):

my thing got cut off at the end

myininaya (myininaya):

+29 is the ending there

myininaya (myininaya):

\[g(f(2))=2^2+9=4+9=13\] i think you can find the last one right?

OpenStudy (anonymous):

thank u..

myininaya (myininaya):

np